100 mL of 0.3N HCl solution were mixed with 200 mL of 0.6 N H2SO4 solution. Calculate the final acidic normality. 

Grade:12th pass

1 Answers

Bhavya
8 years ago
When two acids are mixed:-
N1V1 + N2V2 =N’V’ (final mixture)
100*0.3 + 200*0.6 = N’*300
(100*0.3 + 200*0.6)/300 =N’
0.5 N = N’